### v1

# ESVO2: Direct Visual-Inertial Odometry with Stereo Event Cameras

Abstract
Event-based visual odometry is a specific branch of visual Simultaneous Localization and Mapping (SLAM) techniques, which aims at solving tracking and mapping sub-problems in parallel by exploiting the special working principles of neuromorphic (i.e., event-based) cameras. Due to the motion-dependent nature of event data, explicit data association (i.e., feature matching) under large-baseline view-point changes is hardly established, making direct methods a more rational choice. However, state-of-the-art direct methods are limited by the high computational complexity of the mapping sub-problem and the degeneracy of camera pose tracking in certain degrees of freedom (DoF) in rotation. In this paper, we resolve these issues by building an event-based stereo visual-inertial odometry system on top of our previous direct pipeline Event-based Stereo Visual Odometry [1]. Specifically, to speed up the mapping operation, we propose an efficient strategy for sampling contour points according to the local dynamics of events. The mapping performance is also improved in terms of structure completeness and local smoothness by merging the temporal stereo and static stereo results. To circumvent the degeneracy of camera pose tracking in recovering the pitch and yaw components of general six-DoF motion, we introduce IMU measurements as motion priors via pre-integration. To this end, a compact back-end is proposed for continuously updating the IMU bias and predicting the linear velocity, enabling an accurate motion prediction for camera pose tracking. The resulting system scales well with modern high-resolution event cameras and leads to better global positioning accuracy in large-scale outdoor environments. Extensive evaluations on five publicly available datasets featuring different resolutions and scenarios justify the superior performance of the proposed system against five state-of-the-art methods. We release our pipeline as an open-source software for future research in this field.
Multimedia Material
Supplemental video: https://youtu.be/mdhomp3OCbE
Code: https://github.com/NAIL-HNU/ESVO2.git
I Introduction
Drawing inspiration from biological systems, event cameras represent a revolutionary advancement in sensor technology. These sensors transmit pixel-wise intensity changes, called “events”, asynchronously as they occur [3, 4]. Consequently, they do not produce grayscale images, nor do they adhere to an external clock of fixed capture rate typical of traditional cameras. This asynchronous and differential principle of operation minimizes temporal redundancy, thereby reducing power consumption and bandwidth requirements. Endowed with high temporal resolution and high dynamic range (HDR) capabilities [4], event cameras are qualified to deal with challenging scenarios that are inaccessible to traditional cameras, such as high-speed motion and/or HDR illumination tracking [5, 6, 7, 8, 9, 10, 11, 12], control [13, 14, 15], image synthesis [16, 17, 18, 19] and Simultaneous Localization and Mapping (SLAM) [20, 21, 22, 23, 24, 25, 26, 27].
Although the potential of event cameras is promising, their unconventional output is incompatible with existing modules in visual odometry (VO) / SLAM systems designed for traditional cameras. Event-based VO, on the one hand, inherits the theoretical formulation and parallel design from modern VO solutions. On the other hand, the distinctiveness of event-based methods is witnessed in several aspects, including data processing and representation. Early works using a monocular event camera (i.e., [21, 23]) require a very gentle motion (typically a local-loopy behavior) for the initialization of a local 3D map, based on which the camera pose can be tracked using a 3D-2D registration pipeline. The dependence on dedicated hardware (e.g., a GPU) [21] and the inefficient map expansion [21, 23] restrict their applications on mobile platforms equipped with a limited computation resource and energy supply. More recently, Zhou et al. [1] have proposed an event-based stereo visual odometry (ESVO) pipeline that achieves better accuracy and demonstrates a real-time performance on data collected by a moving agent, such as a flying drone or a ground vehicle. However, this method does not scale well with the streaming rate of event cameras. Real-time performance is only witnessed on event cameras with a small spatial resolution (Q-VGA), e.g., a DAVIS346 (Dynamic and Active-Pixel Vision Sensor) [28]. Also, the failure of tracking in the presence of a sudden variation in angular velocity casts the necessity of introducing an inertial measurement unit (IMU) into the system.
The goal of this paper is to lift the above-mentioned limitations of the original ESVO framework. To this end, we extend ESVO and present a method for visual-inertial odometry (VIO) with a stereo event camera and an IMU. As demonstrated in Fig. LABEL:fig:eye_catcher, the proposed system achieves better mapping and tracking performance than the event-based state of the art (e.g., [1] and [2]) in terms of accuracy and efficiency due to the following efforts.
Contributions:
1.
A novel image-like representation of events, which utilizes their local dynamics for an adaptive accumulation (AA). It is used for efficiently determining pixel locations associated wth instantaneous edges (Sec. IV-A).
2.
An improved solution to the mapping subproblem by leveraging both the temporal-stereo and static-stereo configurations and, additionally, a fast block-matching scheme in the static-stereo operation (Sec. IV-B).
- 3.
4.
An extensive experimental evaluation on five publicly available datasets and comparing against five methods, reporting state-of-the-art performance in accuracy and efficiency (Sec. VI). It also shows that the system scales well with event cameras of different spatial resolutions and is able to run stably in real time on a standard CPU with data of up to VGA resolution. Besides, our dataset and implementation are open-sourced.
Outline: The rest of the paper is organized as follows. First, a literature review of relevant work is provided in Sec. II. Second, the proposed framework is then overviewed from a systematic perspective (Sec. III). Then, we discuss our method by detailing each item listed in the contribution (Secs. IV and V). Finally, experimental results are provided in Sec. VI, followed by the conclusion (Sec. VII).
II Related Work
Like its standard-vision counterparts, event-based VO/SLAM also aims to solve simultaneously the mapping and tracking sub-problems in a recursive manner. The main challenge therein is to answer the following question: How to process raw events in a way that photometric and geometric constraints can be established to infer the camera pose and scene depth information? To this end, several data processing pipelines that utilize different sensor configurations and exploit various data representations have been proposed, as summarized in Tab. I. From the perspective of how event data are processed, existing methods (also including event-based VIO) can be mainly divided into two categories (2nd column of Tab. I): indirect methods and direct methods.
Indirect Methods: To build on top of existing indirect methods (e.g., [29, 30]) using standard cameras, researchers have resorted to developing hand-crafted features from event data, such as event corners [31, 32, 33, 34], which are typically adapted from the original Harris [35] and FAST [36] methods. Additionally, strategies for tracking event corners are presented by [33, 37]. Such event features enable straightforward application of mature geometric tools, such as epipolar geometry [38, 39] and Perspective-n-Point (PnP) methods [40, 41]. Despite the success of these feature-based solutions (e.g., [42] and [43]), event features are not as theoretically robust as their standard-vision counterparts. This is due to the motion-dependent nature of event data, which sometimes leads to incomplete observation of junctions. Consequently, feature matching can easily fail in a sudden variation of the event camera’s motion. Another mainstream strategy for feature detection and tracking is inspired by the motion compensation method [44], a unified pipeline for event-based model fitting. This strategy is widely witnessed in event-based VIO pipelines [45, 24, 2], which typically build features from motion-compensated event sets [7] or event histograms [46] and further fuse with inertial measurements via either a Kalman filter [47] or keyframe-based nonlinear optimization [48].
Direct Methods: Unlike feature-based methods, direct methods refer to those that directly process either events or raw pixel information via some intermediate representation of event data (4th column of Tab. I). Based on the constant-brightness assumption in logarithmic scale, Kim et al.[21] proposed the first direct method, consisting of three interleaved probabilistic filters to solve the sub-problems of mapping, camera pose tracking and, additionally, recovering the intensity information. To justify that recovering the intensity information is not needed and that a GPU is not required, Rebecq et al.[23] proposed a geometric approach. The mapping module determined the 3D location of scene structures by searching for local maxima of ray densities (represented by a Disparity Space Image – DSI), and the camera pose was estimated by 3D-2D registration, aligning the local 3D map to the 2D event locations on the image plane. These two pioneering works are, however, limited by the requirements of gentle motion in the initialization and slow expansion of the local map.
To overcome these limitations, Zhou et al. [1] presented the first event-based VO pipeline (ESVO) using a stereo event camera. The method exploits spatio-temporal consistency of the events across the image planes of the cameras to solve both localization and mapping sub-problems of VO. Nevertheless, ESVO does not achieve real-time performance for camera spatial resolutions of pixels (VGA) or higher. This is mainly due to the large number of redundant operations in the mapping module, which are originally caused by the way contour points are determined. Besides, we observe that ESVO’s tracking module sometimes cannot fully recover the yaw and pitch components in general 6-DoF motion, which degrades the accuracy of the recovered trajectory. To resolve these issues, we recently proposed an IMU-aided version [49] of ESVO. By employing an efficient strategy for sampling contour points, it scales better with modern event cameras of VGA resolution. The degeneracy issue of camera pose tracking is circumvented by introducing as a prior the gyroscope measurements via pre-integration. We also noticed the recent emergence of a direct method using a stereo event camera and an IMU [50]. It is built on top of ESVO, with an additional back-end that refines camera poses using IMU pre-integration constraints. Although it outperforms ESVO in terms of trajectory accuracy, the core limitations of ESVO mentioned above are not well resolved.
Method (In)Direct Sensors Event representation Real-time performance Evaluation datasets Kim et al. [21] Direct Monocular Raw events (128 128 px) on GPU – Rebecq et al. [23] Direct Monocular Raw events & Naive Accum. (240 180 px) on CPU ECD Zhu et al. [45] Indirect Monocular + IMU Raw events ✗ ECD Rebecq et al. [24] Indirect Monocular + IMU Motion-compensated Event Images (240 180 px) on CPU – ESVO [1] Direct Stereo Time Surfaces (346 260 px) on CPU RPG, MVSEC, HKUST Hadviger et al. [42] Indirect Stereo Time Surfaces (346 260 px) on CPU MVSEC, DSEC ESIO [2] Indirect Stereo + IMU Time Surfaces (346 260 px) on CPU HKU, MVSEC, VECtor Wang et al. [43] Indirect Stereo Time Surfaces & Naive Accum. ✗ MVSEC ESVIO [50] Direct Stereo + IMU Time Surfaces (346 260 px) on CPU RPG, MVSEC, ESIM, DSEC Elmoudni et al. [51] Direct Stereo Raw events & Time Surfaces ✗ DSEC Shiba et al. [52] Direct Monocular / Stereo Raw events ✗ RPG, DSEC ES-PTAM [53] Direct Stereo Raw events & Naive Accum. (346 260 px) on CPU RPG, MVSEC, DSEC, EVIMO2, TUM-VIE Niu et al. [49] Direct Stereo + IMU Time Surfaces & Adaptive Accum. (346 260 px) on CPU RPG, DSEC This work Direct Stereo + IMU Time Surfaces & Adaptive Accum. (640 480 px) on CPU RPG, MVSEC, DSEC, VECtor, TUM-VIE
Some other follow-up works of ESVO, such as [51], specifically focus on improving the accuracy of the mapping sub-problem using stereo-fused ray densities [54, 53]. The resulting semi-dense depth maps are sharper and more accurate, however, real-time performance on VGA spatial resolution cameras has never been reported. Finally, Shiba et al.[52] recently extended the Contrast Maximization framework [44, 55] to estimate optical flow and ego-motion from event data. With special consideration on the space-time properties of event data, it can handle overfitting in the event-alignment problem, thus leading to better performance in geometric-model-fitting tasks, including ego-motion estimation. Nevertheless, this theoretical work does not demonstrate real-time performance.
The system proposed in this paper is built on top of our recent ICRA work [49] (also in Tab. I) with the following extensions and improvements:
1.
We improve the AA scheme by incorporating the approach from [56], allowing it to adapt to the variation of event rate without manually adjusting parameters.
2.
We significantly reduce the computational cost in the stereo matching by accelerating the stereo similarity evaluation, enabling the mapping module to run stably at 20 Hz on event data of VGA spatial resolution.
3.
We introduce a back-end using as input the IMU data and past poses to estimate IMU biases and the camera’s velocity, which are then used to provide initial values for the spatio-temporal registration, leading to more efficient and accurate performance of camera pose tracking.
4.
The proposed pipeline is more extensively evaluated, and the performance in terms of efficiency and accuracy is further improved.
III System Overview
Using the original ESVO [1] pipeline as a backbone, we extend it into a stereo VIO system. Given as input the raw events from a calibrated stereo event camera and the inertial measurements from an IMU, the goal is identical to that of ESVO, namely estimating the stereo event camera’s pose in real time and recovering a semi-dense depth map and, furthermore, aiming at a better performance in terms of accuracy and efficiency. Figure -1 displays the flowchart of the system, with the newly added functions highlighted in bold. The entire system can be primarily divided into four modules: pre-processing of event data, mapping, camera-pose estimation, and a back-end that keeps updating the IMU bias and the camera’s linear velocity. The first three modules constitute what is referred to as the front-end.
To keep this section self-contained, let us briefly introduce the functions in each module, with a particular emphasis on new components added to the original ESVO pipeline.
First of all, the pre-processing module generates two event representations, including the time surface (TS) map (refer to details in Sec. III. A of [1]) and our novel representation called the adaptive accumulation (AA) map of events (see Sec. IV-A). Both of them can be refreshed in either an asynchronous manner (e.g., at the occurrence of a certain number of events) or at a fixed rate (e.g., 100 Hz in our implementation) and are stored separately in a database for access by other modules.
Secondly, IMU pre-integration is used as the motion prior to initialize the camera pose estimator, which is formulated as a 3D-2D spatio-temporal registration problem (alignment of the 3D local map projected onto the time-surface domain). The resulting pose estimates are stored in a database of coordinate transforms (e.g., TF in ROS [57]).
Thirdly, the mapping module takes the contour points sampled on the AA maps, time surfaces, and camera pose estimates to keep updating a semi-dense local depth map. Specifically, two kinds of stereo methods are deployed to handle points with different observability properties on the spatio-temporal profile of events. The resulting depth estimates are merged to refresh the local map, which is constantly pushed into a database of global point cloud for potential visualization. Finally, the back-end realizes a sliding-window optimization that keeps updating the estimates of IMU biases and camera velocity at a fixed rate.
The bootstrap of our system creates an initial local map and initializes the unknowns in the back-end (i.e., the IMU biases and camera velocity). The former operation is identical to that in ESVO. A loosely-coupled method is performed in the latter operation, which determines the back-end parameters by aligning two trajectories obtained independently from events and IMU measurements. Different from what has been proposed in the initialization of [50], we find that the refinement of scale is unnecessary as long as the extrinsic parameters are sufficiently accurate.
IV Front-End
We detail our front-end in this section. First, we introduce a novel image-like representation of event data that facilitates efficient sampling of contour points (Sec. IV-A). Second, we discuss our modifications to ESVO’s mapping that lead to improvements in terms of structure completeness and local smoothness (Sec. IV-B). Then, we demonstrate how to further enhance the camera pose estimation by incorporating the IMU pre-integration as motion priors and employing offset-free time surfaces in the spatio-temporal registration (Sec. IV-C).
The front-end interacts with a compact and efficient back-end (Sec. V) that helps suppress drift in the estimated IMU biases, which is key to globally consistent trajectory estimates.
IV-A Pre-processing: Adaptive Accumulation of Events
The computational efficiency of the mapping method in ESVO is limited by several aspects, and one of them is the way that the edge-pixel locations are determined. In ESVO, an edge map in a virtual reference frame is created by applying motion compensation to events occurring within a short time interval (e.g., 10 ms). This operation will become computationally expensive as a pre-processing step when the streaming rate of events exceeds a certain range111The event streaming rate is in proportion to several aspects, including the scene dynamics, scene texture, and spatial resolution of event cameras.. Besides, we observe that the extracted events are typically concentrated in certain regions with large optical flow. To alleviate such an uneven distribution of contour points, the method has to sample a large number of points, which is redundant and becomes a computational burden to mapping. Therefore, a more efficient way is needed to determine contour points.
The synthesized event map obtained by a naive accumulation of events in [23] can be used as an approximate edge map. However, this approximation can become considerably inaccurate (i.e., blurred edges or invisible edges) when the depth varies significantly in the scene. This is because a global threshold for event accumulation cannot handle diverse local dynamics of events. To address this, we propose AA, a novel representation of event data, that can control the amount of events to be accumulated according to the local event dynamics. In [56], the event dynamic is represented by event activity , which is defined as the number of events within a given time interval . A temporal decay strategy is applied to the event activity, which is updated every time an event arrives. The event activity at the arrival of the event can be represented as
where is the decay rate that varies with event activity and is determined by the time interval between two consecutive events:
The method in [56] utilizes the event activity within a time interval to represent the global event dynamics, controlling the decay rate depending on the event activity to generate adaptive-decay time surfaces. Inspired by [56], our method monitors the event activity to determine the termination time of event accumulation. Intuitively, a greater event activity corresponds to higher event dynamics, indicating a shorter duration for event accumulation. We also observe that the event dynamics remain relatively stable within a short time interval (e.g., 10 ms), which leads to event activities converging towards a constant value incrementally, as shown in Fig. 0(a). Assuming the events are generated at equal time intervals (e.g., s), the converged event activity can be roughly calculated as
As seen in Fig. 0(b), this estimation of converged event activity also holds approximately when the time interval is not constant. Therefore, we terminate event accumulation when the event activity converges to obtain a sufficiently sharp image. To handle the spatial variation of local event dynamics, we divide the image plane into small regions and accumulate events independently according to the local event activity.
The generation of an AA map is summarized in Algorithm 1. Specifically, we first divide the image plane evenly into small regions, and the accumulation of events is carried out in each region independently. For each region, the converged event activity is estimated using (3), and then all spatially involved events are sequentially accumulated and used to update the event activity until it converges. The pixel value in an AA map represents the number of events accumulated at this pixel. Pixels with higher values are more likely to be associated with contour points.
To assess the effectiveness of the proposed AA method, we compare it against three image-like representations, including time surfaces [58] and two speed-invariant ones, namely Speed Invariant Learned Corners (SILC) [59] and Threshold-Ordinal Surface (TOS) [60]. As illustrated in Fig. 1, the AA map preserves relatively complete edges with fewest redundant points while keeping the highest signal-to-noise ratio (SNR)222By “SNR”, we mean the ratio between edge points and non-edge points in all the sampled points.. This superior performance can also be justified quantitatively by comparing each representation against the edge map obtained from the corresponding gray-scale image. As shown in Tab. II, the proposed AA representation leads to the highest score using the multi-component structure similarity index measure (3-SSIM) [61] as evaluation metric.
| Event map | TS [62] | SILC [59] | TOS [60] | AA |
| 3-SSIM | 0.1801 | 0.1369 | 0.15203 | 0.3387 |
Contour points used by the following mapping operation are further sampled from each region independently. In general, the higher the AA pixel value in each region, the greater the probability that the pixel will be selected. We compare the sampling result with contour points obtained from ESVO. As shown in Fig. 2, our sampled pixels better capture the edge structures, indicating that our sampling method can cover the edge patterns in the environment using fewer redundant input points.
IV-B Mapping: Depth Estimation
The mapping module aims at recovering depth information and maintaining a local 3D map. In the original ESVO pipeline, these are achieved by constantly propagating recent depth estimates of sampled events to a virtual reference frame, on which overlapped estimates are further fused based on their probabilistic characteristics. As for estimating the depth of each individual event, a two-step method (i.e., a block matching operation followed by a nonlinear refinement) is proposed. In general, the applied static stereo configuration leads to unreliable depth estimation of structures parallel to the stereo baseline. Moreover, the two-step method for depth estimation does not scale well with the number of individual queries. To circumvent these two problems, we propose two modifications: 1) an additional temporal-stereo matching method and 2) a faster design of block matching. We detail each modification in the following.
IV-B1 Temporal Stereo Matching
The static stereo method in [1] can hardly recover accurate depth of structures that are parallel to the baseline of the stereo camera. This is because the spatio-temporal profile of these structures is not distinctive along the baseline, and thus, multiple false-positive matches will be obtained during the block matching operation. Inspired by [63], we introduce the temporal stereo method (i.e., establishing stereo associations between adjacent views of the same camera) to resolve this problem. It is known that as long as the stereo camera does not move in the direction of the baseline, epipolar lines defined between a temporal-stereo pair are no longer parallel to the static-stereo baseline. This assumption always holds for forward moving stereo cameras, e.g., those employed in driving scenarios.
Let’s consider the commonly used horizontal stereo configuration. We divide the sampled contour points from Sec. IV-A into two groups according to their gradient direction on the TS. In the first group, we collect pixels at which the ratio () between the vertical gradient and the horizontal gradient is smaller than a threshold. These pixels are fed to the static stereo method. The remaining sampled pixels go to the second group and are fed to the proposed temporal stereo method.
The key to the event-based temporal-stereo problem is to effectively exploit appearance similarity in event data. This requires the applied event representation, on which the stereo data association is established, to possess the speed-invariant property to some extent. Therefore, the TS used in ESVO [1] is not applicable. To address this, we investigate the temporal stereo matching performance on three representations, namely TOS [60], SILC [59], and our AA. We find that AA is the optimal choice because of its highest contour points ratio. Consequently, the epipolar matching is carried out on successive AA maps of the left camera.
IV-B2 Fast Block Matching
The static stereo method in ESVO [1] is initialized by a block-matching operation, which uses the zero-mean normalized cross-correlation (ZNCC) as a similarity measure:
where denotes the intensity covariance between the patch centered at on image and the patch centered at on image ; and refer to the variance of and , respectively.
| Algorithm | Traditional ZNCC | Fast ZNCC |
| Average time | 0.0317 ms | 0.0057 ms |
Although the original ZNCC leads to satisfactory accuracy, it is notably time-consuming. To speed up, we propose a fast version of ZNCC based on the approach presented in [64]. It adopts an recursive method that calculates the required sums and squared sums for the computation of ZNCC:
where and denote the number of rows and columns of and , the mean intensity of , the sum of the intensities in , and the sum of the squared intensities in . In this way, the approach effectively utilizes intermediate results from previous computations to reduce redundant computation in ZNCC. As shown in Table. III, this recursive method significantly reduces the computational complexity of block matching, leading to a speedup of approximately an order of magnitude.
IV-B3 Depth Merge
Given as input the relative pose between successive AA maps of the left camera, the proposed temporal-stereo method returns the depth information of structures lacking in the static-stereo method. Ideally, these stereo results can be fused in the same way as done in [1], because the temporal-stereo residuals evaluated on AAs and the static-stereo residuals evaluated on TSs approximately obey the Student’s t distribution
where , , are the model parameters, namely the mean, scale, and degree of freedom. However, this is not sensible because the uncertainty of the temporal-stereo results is always much smaller, leading to a bias towards the temporal-stereo estimates. Hence, we simply merge the results of the two stereo methods to obtain a more complete depth map, as shown in Sec. VI-B. Also note that we get rid of the nonlinear refinement of depth in the static-stereo and temporal-stereo methods in our implementation. We find that the absence of the nonlinear refinement guarantees the real-time performance of mapping, while not bringing any notable degradation in the accuracy of camera pose tracking. The effect of omitting the nonlinear refinement is discussed in Sec. VI-E3.
IV-C Localization: Camera Pose Estimation
Our front-end basically follows the tracking module in ESVO [1], which takes as input a TS and a local 3D map, and computes the pose of the stereo rig with respect to the map. Let represent a set of pixels with inverse depth values in the reference frame, and and be the TS and negative TS at time . The objective is, identical to [1], to determine the optimal motion parameters by solving
where the warp function denotes the transformation from local depth points to the latest negative TS image. Besides, denotes the motion parameters, where are the Cayley parameters [65] for rotation and represents the translation.
The spatio-temporal registration module leverages the distance-transform property of time surfaces and formulates the camera pose estimation as a 3D-2D registration problem. The employment of time surfaces smoothed with a blur kernel, on the one hand, enhances the differentiability of the objective function. On the other hand, the true location of edge points is shifted on the image plane, leading to a bias in the registration result. The nonlinear property of the resulting 3D-2D registration problem must be initialized properly, especially in the presence of a sudden variation in the orientation. Also, it may come across a degeneration in recovering the pitch and yaw components, and thus, an additional sensor (e.g., an IMU) is needed as a supplement. We discuss our solutions to these problems in the following.
IV-C1 Offset-free Smoothed Time Surface (OS-TS)
The original TS provides a truncated distance field with slopes ascending towards the true location of contours / edges. The lack of gradients in the truncated area, as shown in Fig. 3(a), leads to a one-sided convergence basin. To this end, a blurring filter is applied to the TS in ESVO [1] to obtain a two-sided convergence basin, as illustrated in Fig. 3(b). The smoothing operation, however, shifts the basin, leading to a notable offset in the location of the edges. This phenomenon can be seen by comparing the region with the highest pixel values in Fig. 3(a) and Fig. 3(b). Consequently, the spatio-temporal registration with blurred TS is no longer optimal because the 3D map is aligned with the shifted observation of edges.
To address this issue, we propose an alternative design, called offset-free smoothed time surface (OS-TS), as outlined in Alg. 2. Specifically, it fills the side lacking valid gradient information in the original TS (Fig. 3(a)) using the pixel values in the corresponding area of the smoothed TS (Fig. 3(b)). As illustrated in Fig. 3(c), the result exhibits no offset while preserving valid gradient information on both sides of the edge. Finally, the negative OS-TS is computed and used in the spatio-temporal registration.
IV-C2 Using IMU Pre-integration as Motion Priors
To deal with sudden variations in orientation and provide a proper initialized value for the spatio-temporal registration, we use as motion priors the pre-integration of an IMU [66]. For an IMU with a 3-axis accelerometer and gyroscope, the measurements of acceleration and angular velocity can be denoted as
where , , and are the bias and noise of the accelerometer and gyroscope in the IMU frame, respectively. The relative translation, velocity, and rotation between two successive pose estimates, e.g., from time to , expressed in the IMU frame , can be calculated by
where denotes the gravity in the IMU coordinate system at time , , and the skew-symmetric matrix of a vector. The linear velocity and biases (i.e., , and ) are initialized empirically and constantly updated in the back-end (see V). The relative velocity and rotation from IMU frame to , namely and , are used as motion priors in the spatio-temporal registration.
V Back-end Optimization
To suppress the drift in the motion estimates, a bundle adjustment (BA) optimization is typically needed as a back-end. One of the most widely used strategies is the sliding-window approach, which jointly refines the poses of all involved key frames and, optionally, the depth information of co-visible 3D points by minimizing the inconsistency among heterogeneous measurements. However, our geometric method lacks explicit data association in the spatio-temporal domain, and thus, such a relaxation-and-refinement scheme [67] is not applicable. Moreover, we observe that the presence of camera poses among the optimization variables does not improve the camera tracking performance. This is because the local map (i.e., the involved point cloud within the window) is a fusion result relying on existing camera poses. It is by no means to establish additional constraints by introducing the 3D point cloud into the sliding-window optimization without relaxation. Therefore, we attempt to create a compact and efficient back-end, which considers only the linear velocity and IMU’s bias as the optimizing variables. Experiments show that the back-end built on this strategy significantly improves the tracking accuracy of our system. (Sec VI-E)
Specifically, we apply a sliding window (shown in Fig. 4), in which the last pose estimates333Note that the involved poses are only used for providing the relative pose constraints rather than being updated. and the corresponding IMU’s measurements are involved. The objective is to refine the estimated linear velocity and IMU biases at the corresponding time of each involved pose so that the relative poses comply maximally with the IMU pre-integration results:
where denotes the optimizing variables (i.e., the linear velocities and the biases of IMU), the involved pose estimates, and the IMU pre-integrated measurements. In addition, denotes the set of IMU pre-integration results in the sliding window. The relative pose residual is specifically defined as:
where and denote the position and rotation, respectively. extracts the vector portion of the quaternion for representing the error state. In detail, the size of the sliding window is set to 5, and we apply the Levenberg-Marquardt (LM) algorithm to solve the optimization problem. The sliding window shifts forward when a new pose estimate is obtained, while the oldest pose estimate and corresponding IMU measurements are removed.
VI Experiments
In this section, we comprehensively evaluate our system. First, we introduce all datasets used and explain why some of them are not applicable for the evaluation (Sec. VI-A). Second, we conduct both qualitative and quantitative comparisons of the proposed mapping module against that of ESVO, justifying the benefits brought by our modifications (Sec. VI-B). Third, we evaluate the overall performance of our system by comparing the localization results against those of various stereo event-based pipelines and demonstrate our superior performance in terms of trajectory accuracy, especially in driving scenarios (Sec. VI-C). Besides, an extended evaluation using our dataset is provided to justify the versatility of our system (Sec. VI-D). Moreover, we investigate the benefits brought by the newly incorporated functions and modules on the tracking performance (Sec. VI-E) and provide a computational-complexity analysis of each function in the entire system (Sec. VI-F). Finally, we discuss some limitations of our method in small-scale and narrow indoor scenes (Sec. VI-G).
VI-A Public Datasets Used
Our system is tested on five publicly available datasets with stereo event cameras and IMUs, which feature different event-camera types, spatial resolutions, and scene geometries, including DSEC [68], VECtor [69], MVSEC [70], rpg [71] and TUM-VIE [72]. DSEC is a large-scale dataset capturing outdoor driving scenes, and the others feature small indoor scenes with different sensor resolutions. Sequences in MVSEC are recorded using a stereo event camera mounted on drones, whereas a hand-held one is used in VECtor, rpg, and TUM-VIE. The characteristics of each dataset are summarized in Tab. IV.
Dataset Cameras Resolution [px] BL [cm] FOV [∘] rpg [71] DAVIS240C 240180 14.7 62.9 MVSEC [70] DAVIS346 346260 10.0 74.8 DSEC [68] Prophesee Gen3.1 640480 59.9 60.1 VECtor [69] Prophesee Gen3 640480 17.0 67.0 TUM-VIE [72] Prophesee Gen4 1280720 11.8 65.0 Ours (Sec. VI-D) DVXplorer 640480 51.0 72.4
Note that some of the other released datasets using stereo event cameras are not applicable to our evaluation. For example, an unsatisfactory rectification is witnessed in the HKU dataset [2] due to inaccuracies in the extrinsic calibration. The ECMD dataset [73] exhibits a significant discrepancy in the streaming rate of events between the left and right event cameras, thus breaking the assumption of temporal coherence between the two cameras. Last but not the least, the M3ED dataset [74] employs event cameras with a spatial resolution of 1280 720 pixels. As discussed in [75], the higher the spatial resolution, the more sensitive each pixel becomes to brightness changes. In urban environments, the excessive number of events recorded in this dataset results in a lack of clear structure information in event images, especially in the presence of trees and green belts. This may result in failures of the proposed direct method.
| Intensity image | ESVO | Our method | |
dsec_city04_a |
dsec_city04_c |
rpg_monitor2 |
rpg_reader |
tumvie_6d |
tumvie_3d |
VI-B Comparison of Mapping: ESVO vs Ours
We compare our mapping results against those of the original ESVO pipeline [1] to demonstrate the performance improvement of our system in mapping. As illustrated in Fig. 5, the results of ESVO exhibit that horizontal structures in the environment are either not reconstructed or inaccurate in depth estimates. This is because the static-stereo method cannot precisely recover the depth of structures that are parallel to the stereo-camera baseline. In comparison, our system additionally incorporates a temporal stereo method, which allows for accurate depth estimation of structures parallel to the stereo-camera baseline. Besides, the sampled points from AA better capture edge structures with less noise (see last row of Fig. 5). All of these lead to better performance in terms of reconstruction completeness and local smoothness.
We also provide a quantitative evaluation, where the figures of merit consist of the mean depth error (defined as ), the median depth error, and the relative error (defined as ), respectively. Sequences that frequently capture horizontal structures are selected for the evaluation. Depth measurements from a LiDAR are adopted as ground truth, and only those within 20 m are used in the evaluation because we find that some of them beyond that range are either less accurate or not aligned well with visual data. As shown in Tab. V, the best results under each metric are highlighted in bold, demonstrating that our method outperforms ESVO in almost all evaluation metrics.
Sequence Depth range Statistic ESVO Ours dsec_city04_a 11.54 m Mean error 1.84 m 1.64 m Median error 1.45 m 1.29 m Relative error 15.15% 12.57% dsec_city04_c 9.75 m Mean error 1.22 m 0.99 m Median error 0.37 m 0.42 m Relative error 11.45% 9.53% dsec_city04_d 11.37 m Mean error 1.72 m 1.59 m Median error 1.83 m 1.35 m Relative error 13.52% 11.20%
VI-C Full System Evaluation
To assess the performance of camera pose tracking of our system, we conduct tests using all five datasets of different resolutions and compare the motion estimation results with another five stereo event-based VO/VIO pipelines [1, 50, 49, 2, 53]. Among them, ESVO [1] is the first stereo event-based VO system that employs a direct method. Successive works, e.g., ESVIO [50] and our ICRA publication [49], are both direct methods that additionally introduce an IMU. Note that a feature-based method, presented in [2], leverages both standard cameras and event cameras. To ensure a fair comparison, only the pure event-based part (denoted by ESIO [2]) is enabled and used as a comparative. Event-based Stereo Parallel Tracking and Mapping (ES-PTAM) [53] is the latest event-only direct stereo VO system, with its mapping module built upon Multi-Camera Event-based Multi-View Stereo (MC-EMVS) [54]. Note that some systems only provide raw data over several sequences. Therefore, we can only use the provided trajectories in our comparison and denote the missing ones with “-” in the tables. We use “failed” to denote cases where the open-sourced systems failed to run successfully.
Dataset Sequence ESVO [1] ES-PTAM [53] ESIO [2] ESVIO [50] ICRA’24 [49] Ours rpg box monitor bin desk reader – – failed failed – – MVSEC indoor_1 indoor_2 – – – – indoor_3 – – failed failed indoor_4 failed failed – – – – failed failed DSEC city04_a 5.59 city04_b city04_c city04_d – – city04_e city04_f city11_a – – city11_b – – – – VECtor robot_normal – – – – corner_slow – – – – hdr_normal – – – – sofa_normal failed failed – – – – failed failed desk_normal – – failed failed – – TUM-VIE 1d_trans failed failed – – 3d_trans failed failed – – 6d_trans failed failed – – failed failed desk failed failed – – desk2 failed failed – –
* “-” represents the lack of experimental results of the algorithm (according to the respective publication).
Dataset Sequence ESVO [1] ES-PTAM [53] ESIO [2] ESVIO [50] ICRA’24 [49] Ours rpg box monitor bin desk reader – – failed failed – – MVSEC indoor_1 indoor_2 – – – – indoor_3 – – failed failed indoor_4 failed failed – – – – failed failed DSEC city04_a city04_b city04_c city04_d – – city04_e city04_f city11_a – – city11_b – – – – VECtor robot_normal – – – – corner_slow – – – – hdr_normal – – – – sofa_normal failed failed – – – – failed failed desk_normal – – failed failed – – TUM-VIE 1d_trans failed failed – – 3d_trans failed failed – – 6d_trans failed failed – – failed failed desk failed failed – – desk2 failed failed – –
To quantitatively evaluate the motion estimation results, we use two standard evaluation metrics: relative pose error (RPE) and absolute trajectory error (ATE) [76]. The results of all methods can be found in Tabs. VII and VII, in which the best results are highlighted in bold. In general, the proposed method (“Ours”) outperforms the others in terms of RPE and ATE. Our method resolves all the remaining issues of ESVO [1], thus leading to more accurate trajectory estimates than ESVO.
In contrast to ESVIO [50], our method advances mainly in the pre-processing and mapping module. The proposed contour-point sampling method and the additionally introduced temporal-stereo operation significantly improve the mapping performance, which benefits the entire system. In addition, our better results justify that having the spatio-temporal-registration residual term incorporated into the back-end by ESVIO [50] does not bring obvious benefits to motion estimation while causing additional computational burden. This may explain why ESVIO [50] does not demonstrate a real-time performance on event data with VGA resolution.
Compared with ESIO [2], the only feature-based method in the table, our direct method makes better use of event data, circumventing the theoretical flaw of event-based feature matching. This advantage of direct methods over feature-based methods is obvious in the results for sequences collected in large-scale environments (e.g., DSEC in Tab. VII). Notably, ESIO cannot operate properly on the TUM-VIE dataset. We believe that one possible reason is that the smoothing effect on the gradient profile of the event representation, caused by the high spatial resolution of the camera, significantly hinders the detection and tracking of ARC* [33], leading to the malfunction of the entire system.
In comparison to [49], our better performance is mainly attributed to the proposed back-end. The better estimated IMU biases and linear velocities at recent poses facilitate camera pose tracking, thus leading to a notable improvement in tracking accuracy. An in-depth discussion on the benefit brought by the back-end is given in Sec. VI-E1. Finally, ES-PTAM [53] performs exceptionally well in small-scale environments and those with repetitive textures, attributing it to the precise depth results via its multi-view stereo mapping method. However, thanks to the effective utilization of IMU data in the back-end, our system significantly outperforms ES-PTAM in large-scale scenarios while being on par in small-scale environments.
Figures 7 and 7 provide a qualitative comparison of the trajectories estimated by ESVO and our method on sequences from rpg, MVSEC, and DSEC data (To deliver a compact and clear visualization, only these trajectories are displayed). Since DSEC does not provide ground-truth trajectories, we use the result of a LiDAR-IMU–based method [77] as ground truth. In short, the trajectories obtained by our method align better with the ground truth, as evidenced by the quantitative analysis.
| Sequence | w/o back-end | w/o OS-TS | w/ refinement | reference | ||||||
| dsec_city04_a | 44.70 (✗) | 3.33 | 0.56 | 32.50 (✓) | ||||||
| dsec_city04_c | 47.55 (✗) | 10.26 | 5.09 | 34.12 (✓) | ||||||
| dsec_city04_d | 47.03 (✗) | 8.87 | 5.04 | 33.45 (✓) | ||||||
| dsec_city04_e | 50.85 (✗) | 3.54 | 0.76 | 30.09 (✓) | ||||||
| dsec_city04_f | 46.41 (✗) | 5.00 | 3.28 | 32.14 (✓) | ||||||
| dsec_city11_a | 49.93 (✗) | 1.85 | 0.487 | 31.96 (✓) | ||||||
| dsec_city11_b | 47.86 (✗) | 10.61 | 4.41 | 36.77 (✓) | ||||||
VI-D Outdoor Evaluation
To further assess the versatility of our system, we collect our own outdoor dataset for evaluation. The employed sensor suite, shown in Fig. 8, consists of a stereo event camera (DVXplorer, 640 480 pixels) with a baseline distance of 51 cm, an embedded 6-axis IMU, and a u-blox ZED-F9P GNSS receiver. The GNSS receiver owns an internal RTK positioning engine that provides ground-truth trajectories for evaluation. All the sensors are calibrated and synchronized in advance.
Our outdoor dataset consists of two subsets (called hnu_mapping and hnu_tracking), which are used for assessing the performance of mapping and localization, respectively. The sequences in the hnu_mapping subset capture scenes that vary in depth range, texture richness, and brightness condition (including HDR scenarios), etc. Some of the mapping results can be found in Fig. 10, which demonstrate good mapping performance as has been witnessed on other publicly available datasets. The hnu_tracking subset consists of two sequences featuring a closed loop (hnu_campus, shown in Fig. LABEL:fig:eye_catcher) and a winding trajectory along a narrow street (hnu_peachlake, shown in Fig. 9), respectively. Compared to the other two stereo event-based visual odometry systems (i.e., ESVO and ESIO), the estimated trajectories of our system align better with the ground truth, especially exhibiting smaller drift in absolute orientation. More importantly, our system is the only one that can operate on these sequences in real time, demonstrating the potential for practical applications.
VI-E Ablation Studies
| Scene | Time Surface (left) | Selected points | Inverse depth map | 3D reconstruction | |
HNU_logo |
HNNU_logo |
HNU_street |
HNU_pavilion |
HNU_statue |
HNU_courtyard |
We conduct several ablation studies to assess the benefits brought by the new functions and modules proposed in this work. Specifically, we use the DSEC dataset [68] and compare the performance of our system (denoted as “reference” in Tab. VIII) with the following features disabled, respectively.
VI-E1 Effect of the Back-end
As said in Sec. V, the presence of back-end facilitates suppressing drift in IMU biases and predicting the linear velocity, thus leading to better performance in camera pose tracking. To justify this, we compare the trajectory results obtained with and without the back-end. In the configuration without the back-end (denoted by “w/o back-end” in Tab. VIII), the gyroscope pre-integration results are still used as an initialization for the spatio-temporal registration, but its bias is not updated. As shown in Tab. VIII, “reference” exhibits notably fewer drift in the estimated trajectories than “w/o back-end”, demonstrating the necessity of constantly updating the belief of the IMU biases and the camera’s linear velocity. With these updated parameters, we obtain better initial values for the spatio-temporal registration and, thus, more accurate motion estimation. Consequently, the improved tracking results lead to more accurate mapping results (especially for the temporal-stereo part) and, in turn, facilitate the tracking module.
It is also worth explaining the difference between the results in Tab. VII (denoted by “ICRA’24”) and those in Tab. VIII. Note that the results obtained with the back-end disabled are less accurate than those reported in [49]. This is because we not only use the gyroscope’s pre-integration results in [49] but also predict the next position using a constant-velocity motion model. In this ablation study, only the rotation pre-integration is performed, which indicates that the constant-velocity motion model can enhance the accuracy of trajectory estimation.
VI-E2 Effect of OS-TS
As claimed in Sec. IV-C, the proposed OS-TS ensures no offset in the location of edges while preserving valid gradient information in the neighboring area. To justify this, we compare the trajectory results obtained with and without OS-TS. In the configuration without OS-TS, the blurred TS (Fig. 3(b)) is used in the 3D-2D spatio-temporal registration. As shown in Tab. VIII, the trajectory errors increase significantly when OS-TS is not employed (denoted by “w/o OS-TS”). This is because the highest pixel values in the blurred TS do not accurately correspond to the positions of the edges in the current scenes. As a result, even when the registration problem is solved successfully, the obtained pose estimation is consistently sub-optimal. Our proposed OS-TS addresses this issue by filling in gradients on the original TS, ensuring the successful resolution of the registration problem and reducing the accumulated error caused by inaccurate pose estimations.
VI-E3 Effect of the Nonlinear Refinement of Depth
In the original ESVO pipeline [1], the disparity estimated by block matching is further refined with a successive nonlinear optimization, aiming at a sub-pixel accuracy. As shown in Tab. VIII, however, we observe that this refinement (denoted by “w/ refinement”) sometimes offers only a marginal improvement in trajectory accuracy compared to the reference method while significantly increasing the runtime of the mapping operation. To guarantee a stable real-time performance at VGA resolution, we choose to omit the nonlinear refinement.
VI-F Computational Efficiency
To assess the computational efficiency, we compare the runtime of ESVO [1], ICRA’24 [49], and the proposed system using a desktop with an Intel Core i7-14700k CPU, as shown in Tab. IX. The tests are performed on the DSEC dataset. All three systems are implemented using hyper-threading technology, and the number of threads occupied by each node is declared inside the parentheses following the node name. The average runtime of each function is profiled, and the numbers in parenthesis (hyper-parameters) denote the amount of data processed by each function. It should be noted that the runtime of our system is primarily influenced by the spatial resolution of the event cameras, proportionally to which the number of points sampled from AA is set. All other hyper-parameters (e.g., the number of points used by each function) are set accordingly, aiming at best performance. As a result, the event streaming rate has minimal influence on our system’s runtime.
[…]
We improve the rendering efficiency of TS and AA by code refactoring, resulting in TS generation being faster than in ESVO and AA generation being faster than in [49]. The efficiency of the front-end is also slightly improved due to better initial values provided by the back-end.
Most of the modifications are made in the mapping module because its performance is mostly relevant to the accuracy and efficiency of the entire system. The improvement in terms of mapping efficiency is due to the following: () The number of points required by the proposed mapping module is significantly reduced compared to ESVO because the input points are no longer raw events but precise contour points extracted by the applied AA and contour-point sampling method; () The proposed fast block matching speeds up the static-stereo operation; () We get rid of the depth refinement because we find that the nonlinear optimization does not bring notable improvements in mapping accuracy when the accurate contour points are fed as input; () The newly added temporal-stereo operation causes a runtime increment (5 ms). Nevertheless, it is worthwhile as it is the key to the improvement of the overall mapping quality. Moreover, the optional regularization operation is no longer used because it is time-consuming and has little effect on the front-end’s performance. The “others” entry in Tab. IX includes general operations in the code that cannot be named as functions, such as data transfer and recycle. The subtotal runtime of the mapping module, shown in Fig. 11, demonstrates that our mapping module can stably operate in real time at 20 Hz on input stereo events of VGA resolution. Last but not the least, our back-end runs in another independent thread, taking 5 ms for each update of the optimization variables in the sliding window.
VI-G Limitations
We observe that our system does not perform well on a small number of sequences, and we attribute this to the inherent limitations of direct methods and the specific characteristics of some datasets. The first category of degenerate cases is typically witnessed in small-scale and narrow spaces, e.g., a textureless corridor (shown in Fig. 12(a) and Fig. 12(b)). In this case, edge patterns either appear parallel to the moving direction of the event camera, resulting in unstable visual observations, or are so close to the event camera that the spatio-temporal profile of event data becomes significantly less distinctive, leading to data association failure in mapping and tracking. This situation exposes more the limitations of the sensor than the limitations of the proposed system. A simple cure is to introduce an additional complementary sensor, e.g., a standard camera, which easily enriches visual information. The second type of degeneration occurs when using low-resolution (e.g., 240 180 pixel) event cameras in cluttered environments of repeated textures. In this case, the streaming rate of events becomes lower than that using a high-resolution event camera under the same dynamics, as justified in Fig. 12(c), thus hindering the determination of contour points via the proposed AA. Furthermore, the presence of widely cluttered, repetitive textures in the scene hinders block matching, thereby worsening depth estimation.
VII Conclusion
We have presented an event-based stereo visual-inertial odometry system on top of our previous work ESVO [1]. It is a direct method that solves the tracking and mapping problems in parallel by leveraging the spatio-temporal coherence in the stereo event data. The goal has been to alleviate ESVO’s high computational complexity in mapping and address its degeneracy in camera pose tracking. To this end, we have additionally introduced an IMU and added several modules to the new pipeline. The newly proposed contour-point sampling method reduces the number of points required by the mapping module, significantly reducing runtime. The temporal-stereo operation acts as a complement to the static stereo, enhancing the mapping quality in terms of structure completeness and local smoothness. The tracking method is still formulated as a 3D-2D registration problem, and its degeneracy is addressed via the usage of IMU pre-integration as motion priors and special considerations for better convergence. To maintain an accurate pre-integration, a compact back-end has been proposed to suppress drift in the estimated IMU biases. Extensive experimental evaluation on five publicly available datasets of different resolutions and our own has demonstrated that our system represents the state of the art and fulfills all the claimed contributions. To the best of our knowledge, the system is the first published work that achieves real-time performance using a standard CPU on event cameras of VGA pixel resolution. The software and datasets used for evaluation are open-sourced.
Acknowledgement
The author would like to thank Mr. Peiyu Chen, Dr. Zhe Liu, and Mr. Suman Ghosh for providing the raw results of [53, 2, 50], respectively, as baselines used in the evaluation. We thank Mr. Jinghang Li and Mr. Kaizhen Sun for the help in the data collection. We also thank Dr. Yi Yu for proof reading. This work was supported by the National Key Research and Development Project of China under Grant 2023YFB4706600.